OpenStudy (anonymous):

Determine the slope of the line that passes through (-11,-8) and (6,16)

OpenStudy (mathstudent55):

To find the slope of a line given two of its points do this: Subtract the y-coordinates. Subtract the x-coordinates. Divide the first subtraction by the second subtraction. Make sure you dot he two subtractions in the same order.

OpenStudy (mathstudent55):

Let's start. Can you subtract the y-coordinates and show what you get?

OpenStudy (anonymous):

\[\frac{ y_2-y_1 }{ x_2-x_1 }\] In your case \[(x_1,y_1)=(-11,-8)\] \[(x_2,y_2)=(6,16)\] Can you fill in the information?

OpenStudy (anonymous):

\[y_2-y_1=16-(-8)=16+8=24\] \[x_2-x_1=6-(-11)=6+11=17\] \[\frac{ 24 }{ 17 }\] It can't be simplified because 17 is a prime number meaning it can only be divided by 1 and itself/